# Transitive Anosov flows on non-compact manifolds

### v1

@@ 3.2. Manifolds with RFRS fundamental groups and their covers
Since $\pi_{1}(M)\curvearrowright P_{\varphi}$ is Anosov-like, $\pi_{1}(\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu)$ as a subgroup of $\pi_{1}(M)$, inherits (A1). Axiom (A3) is vacuously satisfied since $\varphi$ is an Anosov flow and does not have any singular orbit. Finally, we check (A2): By assumption $\mkern 1.5mu\overline{\mkern-1.5mu\varphi\mkern-1.5mu}\mkern 1.5mu$ has a periodic orbit, which is equivalent to having a non-trivial element $h\in\pi_{1}(\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu)$ such that $h\cdot x=x$ for some $x\in\mathcal{P}_{\varphi}$. By item (iii) of Theorem 2.2, $\pi_{1}(M)\cdot\mathcal{F}^{s}(x)$ and $\pi_{1}(M)\cdot\mathcal{F}^{s}(x)$ are both dense in $\mathcal{P}_{\varphi}$. Since $\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu$ is a regular cover, $\pi_{1}(\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu)$ is a normal subgroup of $\pi_{1}(M)$. In particular, for any $g\in\pi_{1}(M)$, the element $ghg^{-1}$ is in $\pi_{1}(\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu)$ and fixes the point $g\cdot x$. Thus, each leaf of $\pi_{1}(M)\cdot\mathcal{F}^{s}(x)$ (resp. $\pi_{1}(M)\cdot\mathcal{F}^{s}(x)$) are fixed by an element of $\pi_{1}(\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu)$, so Axiom (A2) is satisfied. ∎
@@
We first show that $\mkern 1.5mu\overline{\mkern-1.5mu\varphi\mkern-1.5mu}\mkern 1.5mu$ must be transitive when it has a periodic orbit. By Lemma 4.4, $\pi_{1}(\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu)\curvearrowright P_{\varphi}$ satisfies (A1), (A2), and (A3). Since $\varphi$ is $\mathbb{R}$-covered, $P_{\varphi}$ does not have any nonseparated leaves, so (A4) is trivially satisfied. Therefore, $\pi_{1}(\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu)\curvearrowright P_{\varphi}$ is an Anosov-like action. Since $\varphi$ is $\mathbb{R}$-covered, $P_{\varphi}$ is trivial or skew, and Proposition 4.3 implies that $\pi_{1}(\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu)$ is transitive, so $\mkern 1.5mu\overline{\mkern-1.5mu\varphi\mkern-1.5mu}\mkern 1.5mu$ is transitive.
@@
To prove the second part of the statement, one first observes that if $\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu=\widetilde{M}$ then $\mkern 1.5mu\overline{\mkern-1.5mu\varphi\mkern-1.5mu}\mkern 1.5mu$ does not have any periodic orbit, so $\mkern 1.5mu\overline{\mkern-1.5mu\varphi\mkern-1.5mu}\mkern 1.5mu$ is not transitive. Now, suppose that $\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu\neq\widetilde{M}$ and $\mkern 1.5mu\overline{\mkern-1.5mu\varphi\mkern-1.5mu}\mkern 1.5mu$ is not transitive. Then, by the above, $\mkern 1.5mu\overline{\mkern-1.5mu\varphi\mkern-1.5mu}\mkern 1.5mu$ does not have any periodic orbit, i.e., $\pi_{1}(\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu)$ acts freely on $P_{\varphi}$. In particular, $\pi_{1}(\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu)$ also acts freely on the leaf spaces $\Lambda(\mathcal{F}^{s})$ and $\Lambda(\mathcal{F}^{u})$, each of which are homeomorphic to $\mathbb{R}$. By the classical Hölder’s Theorem, we deduce that $\pi_{1}(\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu)$ is abelian, and hence an abelian normal subgroup of $\pi_{1}(M)$. By results of Plante [Pla81] and Ghys [Ghy84] (see, e.g., [BM25, Theorem 2.11.7] for a proof), $\varphi$ is either a lift of the geodesic flow on the unit tangent bundle of some negatively curved orbifold $\Sigma$ or the suspension flow of an Anosov diffeomorphism. We discuss these further in cases.
@@
[…]
@@
In order to apply Proposition 4.3 and the other tools of [BBM24], to deduce the transitivity on regular covers given by Theorem 4.1, it was essential to prove that the action of the subgroup of $\pi_{1}(\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu)$ acted Anosov-like. By Lemma 4.4, the only additional fact that was needed was Axiom (A4), i.e., that branching leaves where fixed. While this is vacuously true for $\mathbb{R}$-covered flows, it is very far from true in general, as we have seen in Theorem 3.8. Thus, to have a generalization of Theorem 4.1 to the $-\mathbb{R}$-covered setting, one needs to further assume that Axiom (A4) is satisfied:
@@
Let $\varphi\colon M\to M$ be a transitive Anosov flow. Let $\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu$ be an infinite regular cover of $M$ and let $\mkern 1.5mu\overline{\mkern-1.5mu\varphi\mkern-1.5mu}\mkern 1.5mu$ be the lift of $\varphi$ to $\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu$. Suppose that all branching leaves in $\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu$ are periodic, and the orbit space $\mathcal{P}_{\varphi}$ does not contain any infinite chain of lozenges that satisfies conditions (ii) and (iii) in Definition 4.2. Then $\mkern 1.5mu\overline{\mkern-1.5mu\varphi\mkern-1.5mu}\mkern 1.5mu$ is transitive.
@@ 5. Non-compact Anosov flows and homotopical characterization of suspensions
Let $\varphi\colon M\to M$ be a suspension Anosov flow. Let $\alpha_{1},\alpha_{2}$ be two distinct periodic orbits of $\varphi$. Do an attracting DA bifurcation on $\alpha_{1}$ and a repelling one on $\alpha_{2}$. Let $V_{1}$, resp. $V_{2}$, be a tubular neighborhood of $\alpha_{1}$, resp. $\alpha_{2}$, with boundary transverse to the DA flow. By [BBY17], we can use an orientation-preserving diffeomorphism $f\colon\partial V_{1}\to\partial V_{2}$ to glue along the boundary tori, obtaining a new (compact) manifold $M_{f}=M/f$, and an Anosov flow $\varphi_{f}$ on $M_{f}$. Since suspension flows are transitive, by [BBY17, Lemma 7.6], $\varphi_{f}$ is also transitive.
@@
[…]
The construction above can be easily extended to build many different examples satisfying the conclusion of the theorem. Note that, given the main result of Fenley in [Fen22], it seems plausible that many (maybe all?) $-\mathbb{R}$-covered Anosov flows on hyperbolic $3$-manifolds have the property that only finitely many periodic orbits are not alone in their free homotopy class.
@@
By construction, $M_{f}$ has a transverse torus $T=\partial V_{1}\sqcup_{f}\partial V_{2}$. The stable foliation and the unstable foliation, when restricted to $T$, have two Reeb components with two closed leaves; see Figure LABEL:fig:_transverse_torus (a). On the orbit space level, projections of lifts of close leaves from the same foliation are nonseparated in their leaf spaces, and projections of lifts of adjacent close leaves from different foliations make perfect fits. The closed leaves on the torus $T$ come from the intersection of $T$ with cylindrical stable and unstable leaves that we denote by $l_{1}^{s},l_{2}^{s}$ and $l_{1}^{u},l_{2}^{u}$ respectively. Lifting to the universal cover, there exists $g\in\pi_{1}(T)<\pi_{1}(M_{f})$ that fixes these leaves $\widetilde{l}^{s}_{1},\widetilde{l}^{s}_{2},\widetilde{l}^{u}_{1},\widetilde{l}^{u}_{2}$, so $g$ fixes points $\widetilde{\alpha}_{1},\widetilde{\alpha}_{2},\widetilde{\beta}_{1},\widetilde{\beta}_{2}$ on each of these leaves in $\mathcal{P}_{\varphi_{f}}$. By Proposition 2.7, these fixed points are corners of a $g$-invariant chain of lozenges, so they correspond to periodic orbits $\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}$ of $\varphi_{f}$ that are freely homotopic to each other (and to $g$ as unoriented curves).
@@
By Proposition 2.7, there exists a chain of lozenge $\mathcal{C}$ containing $\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}$ as corners. Now, suppose that there exists a periodic orbit $\gamma$ of $\varphi_{f}$ such that its free homotopy class $[\gamma]\neq[g]$ and $\gamma$ is not unique in $[\gamma]$. We have the following cases.
@@
Suppose that $\gamma$ intersects $T$ transversely. Then by Proposition 2.7, the projection of $\widetilde{\gamma}$ (the lift of $\gamma$ to $\widetilde{M_{f}}$) to $\mathcal{P}_{\varphi_{f}}$ is a corner of some lozenge. Since $\gamma$ crosses $T$, this corner has to lie in the interior of a lozenge in $\mathcal{C}$. But this is impossible by the non-corner criterion (see [BFM25, Lemma 2.29] or [BM25, Lemma 2.4.7]), which implies that any point in the interior of a lozenge of $\mathcal{C}$ cannot be a corner; see shaded region in Figure LABEL:fig:_transverse_torus (b).
@@
Suppose that $\gamma$ does not cross $T$, i.e., $\gamma\subset M_{f}\setminus T$. Let $\gamma^{\prime}$ be another (possibly the same) orbit in the free homotopy class of $\gamma$. By the first part $\gamma^{\prime}$ is also in $M_{f}\setminus T$. Consider a free homotopy between $\gamma$ and $\gamma^{\prime}$. Either there exists one that stays in $M_{f}\setminus T$ or all of them must cross the torus $T$. In the first case, we would then get a nontrivial free homotopy between orbits of the DA of the suspension flow on $M$, which is absurd. In the second case, we would have that $\gamma$ is freely homotopic to an element in $\pi_{1}(T)$, but by construction, the only periodic orbits of $\varphi_{f}$ that are freely homotopic into $T$ are $\alpha_{1},\alpha_{2},\beta_{1}$ and $\beta_{2}$, which proves the claim.∎
@@
[…]

### v2

@@ 3.2. Manifolds with RFRS fundamental groups and their covers
Since $\pi_{1}(M)\curvearrowright\mathcal{P}_{\varphi}$ is Anosov-like, $\pi_{1}(\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu)$ as a subgroup of $\pi_{1}(M)$, inherits (A1). Axiom (A3) is vacuously satisfied since $\varphi$ is an Anosov flow and does not have any singular orbit. Finally, we check (A2): By assumption $\mkern 1.5mu\overline{\mkern-1.5mu\varphi\mkern-1.5mu}\mkern 1.5mu$ has a periodic orbit, which is equivalent to having a nontrivial element $h\in\pi_{1}(\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu)$ such that $h\cdot x=x$ for some $x\in\mathcal{P}_{\varphi}$. By item (iii) of Theorem 2.2, $\pi_{1}(M)\cdot\mathcal{F}^{s}(x)$ and $\pi_{1}(M)\cdot\mathcal{F}^{s}(x)$ are both dense in $\mathcal{P}_{\varphi}$. Since $\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu$ is a regular cover, $\pi_{1}(\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu)$ is a normal subgroup of $\pi_{1}(M)$. In particular, for any $g\in\pi_{1}(M)$, the element $ghg^{-1}$ is in $\pi_{1}(\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu)$ and fixes the point $g\cdot x$. Thus, each leaf of $\pi_{1}(M)\cdot\mathcal{F}^{s}(x)$ (resp. $\pi_{1}(M)\cdot\mathcal{F}^{s}(x)$) are fixed by an element of $\pi_{1}(\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu)$, so Axiom (A2) is satisfied. ∎
@@
We first show that $\mkern 1.5mu\overline{\mkern-1.5mu\varphi\mkern-1.5mu}\mkern 1.5mu$ must be transitive when it has a periodic orbit. By Lemma 4.4, $\pi_{1}(\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu)\curvearrowright\mathcal{P}_{\varphi}$ satisfies (A1), (A2), and (A3). Since $\varphi$ is $\mathbb{R}$-covered, $P_{\varphi}$ does not have any nonseparated leaves, so (A4) is trivially satisfied. Therefore, $\pi_{1}(\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu)\curvearrowright P_{\varphi}$ is an Anosov-like action. Since $\varphi$ is $\mathbb{R}$-covered, $\mathcal{P}_{\varphi}$ is trivial or skew, and Proposition 4.3 implies that $\pi_{1}(\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu)$ is transitive, so $\mkern 1.5mu\overline{\mkern-1.5mu\varphi\mkern-1.5mu}\mkern 1.5mu$ is transitive.
@@
To prove the second part of the statement, one first observes that if $\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu=\widetilde{M}$ then $\mkern 1.5mu\overline{\mkern-1.5mu\varphi\mkern-1.5mu}\mkern 1.5mu$ does not have any periodic orbit, so $\mkern 1.5mu\overline{\mkern-1.5mu\varphi\mkern-1.5mu}\mkern 1.5mu$ is not transitive. Now, suppose that $\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu\neq\widetilde{M}$ and $\mkern 1.5mu\overline{\mkern-1.5mu\varphi\mkern-1.5mu}\mkern 1.5mu$ is not transitive. Then, by the above, $\mkern 1.5mu\overline{\mkern-1.5mu\varphi\mkern-1.5mu}\mkern 1.5mu$ does not have any periodic orbit, i.e., $\pi_{1}(\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu)$ acts freely on $\mathcal{P}_{\varphi}$. In particular, $\pi_{1}(\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu)$ also acts freely on the leaf spaces $\Lambda(\mathcal{F}^{s})$ and $\Lambda(\mathcal{F}^{u})$, each of which are homeomorphic to $\mathbb{R}$. By the classical Hölder’s Theorem, we deduce that $\pi_{1}(\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu)$ is abelian, and hence an abelian normal subgroup of $\pi_{1}(M)$. By results of Plante [Pla81] and Ghys [Ghy84] (see, e.g., [BM25, Theorem 2.11.7] for a proof), $\varphi$ is either a lift of the geodesic flow on the unit tangent bundle of some negatively curved orbifold $\Sigma$ or the suspension flow of an Anosov diffeomorphism. We discuss these further in cases.
@@
Suppose that $\varphi$ is a lift of the geodesic flow on $T^{1}\Sigma$, so the orbit space of $\varphi$ is skew. Let $h\colon(x,y)\mapsto(x+1,y+1)$ be the one-step-up map on the skew plane. Then by Theorem 2.11.6 in [BM25], $\pi_{1}(\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu)$ is contained in the cyclic center $\langle h^{k}\rangle$ of $\pi_{1}(M)$. The cover associated to $\langle h\rangle$ is $T^{1}\widetilde{\Sigma}$, and so $\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu$ is obtained as a (self)-cover $T^{1}\widetilde{\Sigma}$ by unwrapping the fibers.
@@
In order to apply Proposition 4.3 and the other tools of [BBM24], to deduce the transitivity on regular covers given by Theorem 4.1, it was essential to prove that the action of the subgroup of $\pi_{1}(\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu)$ acted Anosov-like. By Lemma 4.4, the only additional fact that was needed was Axiom (A4), i.e., that branching leaves where fixed. While this is vacuously true for $\mathbb{R}$-covered flows, it is very far from true in general, as we have seen in Theorem 3.9. Thus, to have a generalization of Theorem 4.1 to the non-$\mathbb{R}$-covered setting, one needs to further assume that Axiom (A4) is satisfied:
@@
Let $\varphi\colon M\to M$ be a transitive Anosov flow. Let $\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu$ be an infinite regular cover of $M$ and let $\mkern 1.5mu\overline{\mkern-1.5mu\varphi\mkern-1.5mu}\mkern 1.5mu$ be the lift of $\varphi$ to $\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu$. Suppose that all branching leaves in $\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu$ are periodic, and the orbit space $\mathcal{P}_{\varphi}$ does not contain any infinite chain of lozenges that satisfies conditions (i) and (ii) in Definition 4.2. Then $\mkern 1.5mu\overline{\mkern-1.5mu\varphi\mkern-1.5mu}\mkern 1.5mu$ is transitive.
@@ 5. Non-compact Anosov flows and homotopical characterization of suspensions
Let $\varphi\colon M\to M$ be a suspension Anosov flow. Let $\alpha_{1},\alpha_{2}$ be two distinct periodic orbits of $\varphi$. Do an attracting DA bifurcation on $\alpha_{1}$ and a repelling one on $\alpha_{2}$. Let $V_{1}$, resp. $V_{2}$, be a tubular neighborhood of $\alpha_{1}$, resp. $\alpha_{2}$, with boundary transverse to the DA flow. By [BBY17], we can use an orientation-preserving diffeomorphism $f\colon\partial V_{1}\to\partial V_{2}$ to glue along the boundary tori, obtaining a new (compact) manifold $M_{f}=M/\penalty 50f$, and an Anosov flow $\varphi_{f}$ on $M_{f}$. Since suspension flows are transitive, by [BBY17, Lemma 7.6], $\varphi_{f}$ is also transitive.
@@
For the flow $\varphi_{f}\colon M_{f}\to M_{f}$ built in Construction 1, there exists a non-compact cover $\mkern 1.5mu\overline{\mkern-1.5muM_{f}\mkern-1.5mu}\mkern 1.5mu$ of $M_{f}$ such that the lifted flow $\mkern 1.5mu\overline{\mkern-1.5mu\varphi_{f}\mkern-1.5mu}\mkern 1.5mu\colon\mkern 1.5mu\overline{\mkern-1.5muM_{f}\mkern-1.5mu}\mkern 1.5mu\to\mkern 1.5mu\overline{\mkern-1.5muM_{f}\mkern-1.5mu}\mkern 1.5mu$ is transitive and every periodic orbit of $\mkern 1.5mu\overline{\mkern-1.5mu\varphi_{f}\mkern-1.5mu}\mkern 1.5mu$ is alone in its free homotopy class.
The construction above can be easily extended to build many different examples satisfying the conclusion of the theorem. Note that, given the main result of Fenley in [Fen22], it seems plausible that many (maybe all?) non-$\mathbb{R}$-covered Anosov flows on hyperbolic $3$-manifolds have the property that only finitely many periodic orbits are not alone in their free homotopy class.
@@
By construction, $M_{f}$ has a transverse torus $\mathbb{T}=\partial V_{1}\sqcup_{f}\partial V_{2}$. The stable foliation and the unstable foliation, when restricted to $T$, have two Reeb components with two closed leaves; see Figure 4(a) (a). On the orbit space level, projections of lifts of closed leaves from the same foliation are nonseparated in their leaf spaces, and projections of lifts of adjacent close leaves from different foliations make perfect fits. The closed leaves on the torus $\mathbb{T}$ come from the intersection of $\mathbb{T}$ with cylindrical stable and unstable leaves that we denote by $l_{1}^{s},l_{2}^{s}$ and $l_{1}^{u},l_{2}^{u}$ respectively. Lifting to the universal cover, there exists $g\in\pi_{1}(\mathbb{T})<\pi_{1}(M_{f})$ that fixes these leaves $\widetilde{l}^{s}_{1},\widetilde{l}^{s}_{2},\widetilde{l}^{u}_{1},\widetilde{l}^{u}_{2}$, so $g$ fixes points $\widetilde{\alpha}_{1},\widetilde{\alpha}_{2},\widetilde{\beta}_{1},\widetilde{\beta}_{2}$ on each of these leaves in $\mathcal{P}_{\varphi_{f}}$. By Proposition 2.7, these fixed points are corners of a $g$-invariant chain of lozenges, so they correspond to periodic orbits $\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}$ of $\varphi_{f}$ that are freely homotopic to each other (and to $g$ as unoriented curves).
@@
By Proposition 2.7, there exists a chain of lozenge $\mathcal{C}$ containing $\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}$ as corners. Now, suppose that there exists a periodic orbit $\gamma$ of $\varphi_{f}$ such that its free homotopy class $\gamma]\neq[g]$ and $\gamma$ is not unique in $\gamma]$. We have the following cases.
(a) Induced foliations on the transverse torus (a); and the projections in the orbit space (b).
@@
Suppose that $\gamma$ intersects $\mathbb{T}$ transversely. Then by Proposition 2.7, the projection of $\widetilde{\gamma}$ (the lift of $\gamma$ to $\widetilde{M_{f}}$) to $\mathcal{P}_{\varphi_{f}}$ is a corner of some lozenge. Since $\gamma$ crosses $\mathbb{T}$, this corner has to lie in the interior of a lozenge in $\mathcal{C}$. But this is impossible by the non-corner criterion (see [BFM25, Lemma 2.29] or [BM25, Lemma 2.4.7]), which implies that any point in the interior of a lozenge of $\mathcal{C}$ cannot be a corner; see shaded region in Figure 4(a) (b).
@@
Suppose that $\gamma$ does not cross $\mathbb{T}$, i.e., $\gamma\subset M_{f}\setminus\mathbb{T}$. Let $\gamma^{\prime}$ be another (possibly the same) orbit in the free homotopy class of $\gamma$. By the first part $\gamma^{\prime}$ is also in $M_{f}\setminus\mathbb{T}$. Consider a free homotopy between $\gamma$ and $\gamma^{\prime}$. Either there exists one that stays in $M_{f}\setminus\mathbb{T}$ or all of them must cross the torus $\mathbb{T}$. In the first case, we would then get a nontrivial free homotopy between orbits of the DA of the suspension flow on $M$, which is absurd. In the second case, we would have that $\gamma$ is freely homotopic to an element in $\pi_{1}(\mathbb{T})$, but by construction, the only periodic orbits of $\varphi_{f}$ that are freely homotopic into $\mathbb{T}$ are $\alpha_{1},\alpha_{2},\beta_{1}$ and $\beta_{2}$, which proves the claim.∎
@@
Let $\mkern 1.5mu\overline{\mkern-1.5mu\varphi\mkern-1.5mu}\mkern 1.5mu_{f}\colon\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu_{f}\to\mkern 1.5mu\overline{\mkern-1.5muM\mkern-1.5mu}\mkern 1.5mu_{f}$ be as in Theorem 5.1. Does $\mkern 1.5mu\overline{\mkern-1.5mu\varphi\mkern-1.5mu}\mkern 1.5mu_{f}$ admit any homologically trivial periodic orbits?